OpenStudy (anonymous):

. Comparing Buying Alternatives. Tammy Monahan is considering the purchase of a home entertainment center. The product attributes she plans to consider and the weights she gives to them are as follows: Portability .1 Sound projection .6 Warranty .3 Tammy related the brands as follows: Portability Sound Projection Warranty Brand A 6 8 7 Brand B 9 6 8 Brand C 5 9 6 Using the consumer buying matrix ( p. 256), conduct a quantitative product evaluation rating for each brad. What other factors is Tammy likely to consider when making her purchase?

OpenStudy (anonymous):

i have no idea what a consumer buying matrix is, but i think you are supposed to compute \[.1\times 6+.6\times 8+.3\times 7\] for the first one others are analogous

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw